OpenStudy (unklerhaukus):

Evaluate in terms of the gamma function \[\int\limits_0^1\sqrt{\frac{1-x^2}{1+x^2}}\text d x\]

OpenStudy (anonymous):

@UnkleRhaukus , where is gamma ??

OpenStudy (blockcolder):

Isn't gamma function the one used for the factorial of non-integer values?

OpenStudy (unklerhaukus):

That is correct http://en.wikipedia.org/wiki/Gamma_function

OpenStudy (unklerhaukus):

In the question though it think we rearrange the equation to a form of the Beta function http://en.wikipedia.org/wiki/Beta_function and then Gamma \[B(m,n)=\frac{\Gamma(n)\Gamma(m)}{\Gamma(m+n)}\]

OpenStudy (blockcolder):

I was thinking of expanding the integrand as a binomial series, but it's too complicated.

OpenStudy (unklerhaukus):

I have the answer if that helps

OpenStudy (anonymous):

i ll just use my calculator lol

OpenStudy (unklerhaukus):

\[=\frac{1}{4}\sqrt \pi\left(\frac{\Gamma (1/4)}{\Gamma(3/4)}-4\frac{\Gamma(3/4)}{\Gamma(1/4)}\right)\]

OpenStudy (blockcolder):

Lemme try my idea. \[\sqrt{\frac{1-x^2}{1+x^2}}=\left ( -1+\frac{2}{1+x^2}\right )^{1/2}\] Okay, I don't get how it became that. =))

OpenStudy (unklerhaukus):

NB: \[\Gamma(\frac1 2)=\sqrt \pi\]

OpenStudy (anonymous):

@UnkleRhaukus , thanks for explaining gamma function. It is a new term which I came across today. Would it be helpful to you if I were to integrate the expression you gave with respect to dx and give you the answer?

OpenStudy (unklerhaukus):

i dont think that is the right method because the definitions of the gamma function and the beta functions are both integrals, but if the answer is particularly illumination i would be interested

OpenStudy (unklerhaukus):

i cant even see how you would be able to integrate it

OpenStudy (experimentx):

let 1+x^2 = u, 2x dx = du x = (u - 1)^(1/2) 1 - x^2 = -u+1 \[ \int_{1}^{0} \frac{(1 - u)^{1/2}}{u^{1/2}2(u-1)^{1/2}} du\]

OpenStudy (experimentx):

bad ... stuck here!!!

OpenStudy (unklerhaukus):

\[x = (u - 1)^{1/2} \]\[x^2=(u-1)\]\[x^2+1=u\] ?\[1 - x^2 ≠ -u+1\]?

OpenStudy (unklerhaukus):

working backwards from the answer \[=\frac{1}{4}B(1/2,1/4)-\frac{1}{2}B(-1/2,3/4)\]

OpenStudy (experimentx):

let x^2 = tan^2A, 2xdx = 2 tanA sec^2AdA dx = sec^2AdA \[\int \sqrt{\cos^2A - \sin^2A} \sec^2 A \;dA = \int (\cos2A)^{1/2} (2\cos2A- 1)^{-1}dA ??? stuck again\]

OpenStudy (unklerhaukus):

wait that looks good, maybe

OpenStudy (unklerhaukus):

\[B(m,n)=2\int\limits_0^{\pi/2} \cos^{(2m-1)}\theta \sin^{(2n-1)}\theta\text d \theta\]

OpenStudy (experimentx):

there is 2 in the middle!!

OpenStudy (unklerhaukus):

\[2\cos2A=1-2\sin^2x\] double angle formula

OpenStudy (unklerhaukus):

i think we have all the pieces

OpenStudy (unklerhaukus):

@experimentX what kind of substitution is this called again i can remember doing this a few years ago \[x^2 = \tan^2A\]